The bottom of a ladder must be placed 3 feet from a wall. The ladder is 12 feet tall. How far up the wall does the ladder reach?


1
Expert's answer
2021-08-31T17:27:18-0400


By the Pythagorean Theorem


"l^2=x^2+y^2"

Given "x=3 ft, l=12ft."


"y=\sqrt{l^2-x^2}=\sqrt{(12)^2-(3)^2}=3\sqrt{15}\ (ft)"

The distance the ladder reaches up the wall is "3\sqrt{15}\approx11.612" feet.